# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v1

@@ Appendix C Evaluating the Efficiency after Node Pruning
Table 5: Evaluation of Efficiency. $\uparrow$ indicates results better than PCST, while $\downarrow$ indicates worse results. (xx%) represents the percentage of tokens relative to PCST.
@@
Non-Retriever 100626.28 OOM -
BM25 [52] 2569.57 17:24 min 0.4394
PCST [34] 2576.66 17:28 min 0.4502
PCST
   w/ $\text{Aligner}_{n_{\text{seed}}=4}$ 496.54 (19.27%) 3:31 min 0.4299$\downarrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=6}$ 698.54 (27.11%) 4:45 min 0.4527$\uparrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=8}$ 905.52 (35.14%) 6:55 min 0.4699$\uparrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=10}$ 1120.75 (43.5%) 8:27 min 0.4785$\uparrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=15}$ 1546.90 (60.04%) 11:50 min 0.4914$\uparrow$
In this section, we evaluate the efficiency improvements brought by node alignment and node pruning in Aligner module. Specifically, we compare different retrieval methods in terms of average token consumption, average inference time, and Hit@1 scores on the test set. The w/ Aligner method is built on the initial retrieval results from PCST, where different seed nodes are selected to perform node pruning. The experiments were conducted on two Nvidia A100 80G GPUs in an inference-only setting. To fully utilize GPU resources and memory, we adjusted the batch size to be as large as possible under different average token conditions. For comparison, the BM25 method retrieves triples and ensures that the token consumption is roughly similar to that of PCST. From Table 5, we observe that when the number of seed nodes ($n_{\text{seed}}$) is set to 6, our Aligner method achieves performance comparable to PCST while utilizing 70% fewer tokens. This demonstrates the method’s ability to effectively prune irrelevant knowledge while preserving the essential information. However, when $n_{\text{seed}}$ is too small (e.g., $n_{\text{seed}}=4$), performance drops significantly, suggesting that some critical information may be pruned during the process. On the other hand, increasing the number of seed nodes to 15 (utilizing 60.04% of the tokens) leads to performance that significantly surpasses PCST. This highlights the strength of our method in efficiently selecting useful knowledge and confirms that, in longer-context scenarios, removing irrelevant information enhances overall results. Additionally, we find that BM25 performs worse than PCST when using a similar number of tokens. This suggests that the PCST retrieval method is better suited for graph databases, as it more effectively captures the connections between knowledge.
Appendix D Impact of Seed Nodes
In this section, we analyze the impact of the number of seed nodes ($n_{\text{seed}}$) on model performance. The experiments are conducted on the WebQSP dataset, where we evaluate the Hit@1 and Accuracy metrics. From the experiments in Figure 3, we observe that the Hit@1 and Accuracy performance peaks when the number of seed nodes is set to 25. Beyond this point (from 25 to 30), the performance starts to decline. This indicates that including too many nodes introduces a significant amount of irrelevant knowledge, which negatively impacts the model. Our pruning strategy effectively eliminates irrelevant knowledge to enhance model performance. On the other hand, when the number of seed nodes is as low as 5, the performance is considerably poor. This suggests that excessive pruning removes crucial knowledge, which is detrimental to performance. This highlights a trade-off: pruning reduces noise and improves performance, but over-pruning leads to the loss of essential knowledge.
Figure 3: Hyperparameters Analysis of the Number of seed nodes.
Appendix E Impact of Align Degree
This section examines how the Align Degree (number of training steps for Aligner module) influences model performance. As shown in Figure 4, we evaluate the Hit@1 and Accuracy metrics on the WebQSP dataset. From the experimental curve of Align Degree, we observe that the Hit@1 and Accuracy metrics peak at around 60 epochs before declining. This indicates that, as training progresses, bridging the representation gap between the graph and language helps the LLM better understand graph data. However, excessive training may lead to overfitting, which disrupts graph information and ultimately causes a drop in performance. This suggests that there is an optimal point in training where the representation alignment is most effective, and training beyond this point can be detrimental.
Figure 4: Hyperparameters Analysis of Align degree.
Appendix F Impact of Top K retrieval
Figure 5 illustrates the impact of varying the Top K retrieval of entities and relations on model performance across Accuracy-score, F1-score, and Hit@1-score. By analyzing the trends in the graphs, we can derive insights into the effect of Top K on model performance. All three metrics show a similar trend: performance improves as K increases, peaks at K = 10, and then declines. This suggests that K = 10 strikes the optimal balance between retrieving relevant results and avoiding noise. Smaller K values may miss relevant information, while larger K values dilute relevance, reducing precision and retrieval quality. These findings highlight the importance of selecting an appropriate K value to maximize performance in retrieval-based systems.
Figure 5: Effect of Top K retrieval.
Appendix G Evaluating the Quality of Summarization Prompts
In this section, we conduct a fine-grained analysis of the quality of summarization for reasoning chain. This analysis is divided into three aspects. First, a case study is used to gain a deeper understanding of the forms of summarization. Second, we evaluate Relevance and Faithfulness to assess whether the generated chain are aligned with the questions and answers. Finally, we compare the QA performance of poorer-quality reasoning chain to highlight their impact.
G.1 Evaluation of Relevance and Faithfulness
To assess the quality of generated reasoning chain in terms of their relevance to question-answer context and their faithfulness to provided contextual information, we adopted LLM-as-a-judge framework [79]. Specifically, we utilized GPT-4o [1] to perform the evaluation. A total of 100 samples were randomly selected for this analysis. The evaluation was conducted using a carefully designed prompt, as illustrated in Figure 6. The results of the experiment are presented in Table 6. The analysis reveals that the summarized reasoning chain exhibit a high degree of relevance and faithfulness. Specifically, the summaries achieved a relevance score of 90%, indicating that the majority of the generated content aligns well with the given question-answer context. A faithfulness score of 87% demonstrates that the summaries adhere largely to the factual information provided by the context, with minimal hallucination.
Figure 6: Evaluation of Relevance and Faithfulness Prompt.
Table 6: Evaluating Relevance and Faithfulness of Reasoning Chain.
Metrics WebQSP
Relevance 90%
Faithfulness 87%
G.2 Quantitative comparison
In this experiment, we evaluate the impact of reasoning quality on QA performance by comparing two approaches: $\text{Reasoning}_{\text{w/ answer}}$, where both the question and answer are included during reasoning chain generation, and $\text{Reasoning}_{\text{w/o answer}}$, where only the question is used for reasoning chain generation. The results in Table 7 demonstrate that incorporating the answer during summary generation ($\text{Reasoning}_{\text{w/ answer}}$) consistently outperforms the approach that excludes the answer ($\text{Reasoning}_{\text{w/o answer}}$) across all evaluation metrics. Specifically, the F1 score of $\text{Reasoning}_{\text{w/ answer}}$ is 0.5445, which is higher than the 0.5209 achieved by $\text{Reasoning}_{\text{w/o answer}}$, indicating that including the answer improves the balance of precision and recall in QA tasks. Similarly, the Hit@1 value increases from 0.7444 to 0.7626, showing that the inclusion of the answer helps the model identify correct nodes more accurately in graph alignment. These findings confirm that using the answer during summary generation leads to higher-quality summaries, which in turn positively impacts QA performance, whereas excluding the answer results in inferior summaries and reduced performance.
Table 7: Quantitative Comparison.
Method WebQSP
@@ G.2 Quantitative comparison
$\text{Reasoning}_{\text{w/ answer}}$ 0.5445 0.7626 0.5700
$\text{Reasoning}_{\text{w/o answer}}$ 0.5209 0.7444 0.5409
G.3 Case study
The goal of this case study is to analyze the effectiveness of summarization in identifying intermediate nodes that are critical for reasoning chains, as shown in 8. By leveraging well-crafted prompts and a graph database, we aim to demonstrate how the summaries enable the logical connection from the query to the final answer. Specifically, we assess the ability of the summarization to extract key intermediate nodes, such as ‘J.K. Rowling’, which play a pivotal role in reasoning. Without summarization, the raw graph database alone is insufficient to provide the answer directly.
@@ G.3 Case study
Table 8: The case for summarizing the reasoning chain. The text from SUMMARIZE_PROMPT, Question, Answer, and Graph DataBase are concatenated as input to generate the Response from LLM. We can see that LLM’s summarization accurately identified middle and critical node in graph, ‘J.K. Rowling’, for answer.
@@
4,the complete idiot’s guide to the world of harry potter (complete idiot’s guide to)
@@
43,harry potter and the order of the phoenix
46,harry potter and the deathly hallows: part i
@@
91,freebase.domain_profile.featured_views,790
91,freebase.domain_profile.featured_views,759
@@
7,book.book.characters,178
43,book.book.characters,16
24,book.book_edition.book,24
@@
Appendix H LLM Prompts
We provide the Summarization Prompt and Generator Prompt in Figure 7 and Figure 8. Summarization prompt is designed to generate a comprehensive summary by combining a given question, its answer, and related graph data. The output should logically connect the question to the answer while incorporating relevant information from the textualized graph data. The generator prompt varies by dataset. For WebQSP and SceneGraphs, it requires answering a question based on textualized graph data. For ExplaGraphs, it determines whether two arguments support or counter each other, providing the answer as either ‘support’ or ‘counter’.
@@ Appendix H LLM Prompts
Figure 7: Prompt for Summarization.
@@
Figure 8: Prompt for Generator.
Appendix I Limitations

### v2

@@ 5.8. Evaluating the Efficiency of Node Pruning
Table 4. Evaluation of Efficiency with and without node pruning. $\uparrow$ indicates results better than G-Retriever, while $\downarrow$ indicates worse results. (xx%) represents the percentage of tokens relative to G-Retriever. OOM means out of memory.
@@
No Pruning Non-Retriever 100626.28 (387.02%) OOM -
BM25 (Robertson et al., 2009) 2569.57 (99.72%) 17:24 min 0.4394
Contriever (Izacard et al., [n. d.]) 2593.28 (100.01%) 17:37 min 0.4429
G-Retriever (Bienstock et al., 1993) 2576.66 (100.00%) 17:28 min 0.4502
Node Pruning with
Different Degree
Align-GRAG
@@
[…]
In this section, we evaluate the efficiency improvements brought by node pruning in Aligner module. Specifically, we compare different retrieval methods in terms of average token consumption, average inference time, and Hit@1 scores on the test set of WebQSP. The w/ Aligner method is built on the initial retrieval results from PCST as the same as G-Retriever (He et al., 2024), where different seed nodes are selected to perform node pruning. The experiments were conducted in an inference-only setting. To fully utilize GPU resources and memory, we adjusted the batch size to be as large as possible under different average token conditions. For comparison, the BM25 and Contriever methods retrieve triples and ensure that the token consumption is roughly similar to that of G-Retriever. From Table 4, we observe that when the number of seed nodes ($n_{\text{seed}}$) is set to 6, our Aligner method achieves performance comparable to G-Retriever while only utilizing 27.11% tokens. This demonstrates the method’s ability to effectively prune irrelevant knowledge while preserving the essential information. However, when $n_{\text{seed}}$ is too small (e.g., $n_{\text{seed}}=4$), performance drops significantly, suggesting that some critical information may be pruned during the process. On the other hand, increasing the number of seed nodes to 15 (utilizing 60.04% of the tokens) leads to performance that significantly surpasses G-Retriever. This highlights the strength of our method in efficiently selecting useful knowledge and confirms that, in longer-context scenarios, removing irrelevant information enhances overall results. Additionally, we find that BM25 and Contriever both perform worse than G-Retriever when using a similar number of tokens. This suggests that the PCST retrieval method is better suited for graph databases, as it more effectively captures the connections between knowledge.
5.9. Impact of Hyperparameters
In this section, we analyze how hyperparameters affect performance, including: the Number of Seed Nodes $n_{\text{seed}}$, Align Degree (training step) and Top K retrieval.
5.9.1. Impact of Seed Nodes
We analyze the impact of the number of seed nodes ($n_{\text{seed}}$) on model performance. The experiments are conducted on the WebQSP dataset, where we evaluate the Hit@1, F1 and Accuracy metrics. From the experiments in Figure 5, we observe that the Hit@1, F1 and Accuracy performance peaks when the number of seed nodes is set to 25. Beyond this point (from 25 to 30), the performance starts to decline. This indicates that including too many nodes introduces a significant amount of irrelevant knowledge, which negatively impacts the model. Our pruning strategy effectively eliminates irrelevant knowledge to enhance model performance. On the other hand, when the number of seed nodes is as low as 5, the performance is considerably poor. This suggests that excessive pruning removes crucial knowledge, which is detrimental to performance. This highlights a trade-off: pruning reduces noise and improves performance, but over-pruning leads to the loss of essential knowledge.
Figure 5. Hyperparameters Analysis of the Number of seed nodes.
5.9.2. Impact of Align Degree
This section examines how the Align Degree (number of training steps for Aligner module) influences model performance. As shown in Figure 6, we evaluate the Hit@1, F1 and Accuracy metrics on the WebQSP dataset. From the experimental curve of Align Degree, we observe that the Hit@1, F1 and Accuracy metrics peak at around 60 epochs before declining. This indicates that, as training progresses, bridging the representation gap between the graph and language helps the LLM better understand graph data. However, excessive training may lead to overfitting, which disrupts graph information and ultimately causes a drop in performance. This suggests that there is an optimal point in training where the representation alignment is most effective, and training beyond this point can be detrimental.
Figure 6. Hyperparameters Analysis of Align degree.
5.9.3. Impact of Top K retrieval
Figure 7 illustrates the impact of varying the Top K retrieval of entities and relations on model performance across Hit@1, F1 and Accuracy. By analyzing the trends in the graphs, we can derive insights into the effect of Top K on model performance. All three metrics show a similar trend: performance improves as K increases, peaks at K = 10, and then declines. This suggests that K = 10 strikes the optimal balance between retrieving relevant results and avoiding noise. Smaller K values may miss relevant information, while larger K values dilute relevance, reducing precision and retrieval quality. These findings highlight the importance of selecting an appropriate K value to maximize performance in retrieval-based systems.
Figure 7. Effect of Top K retrieval.
5.10. Evaluating the Quality of Reasoning Chain
In this section, we conduct a fine-grained analysis of the quality of the reasoning chain by LLM summarization. This analysis is divided into two aspects. First, we evaluate Relevance and Faithfulness to assess whether the generated chains are aligned with the questions and answers. Second, we compare the QA performance of the naive-quality reasoning chain to highlight our method’s impact. What is more, a case study is used to gain a deeper understanding of the forms of summarization in Section 5.11.
5.10.1. Evaluation of Relevance and Faithfulness
To assess the quality of generated reasoning chain in terms of their relevance to question-answer context and their faithfulness to provided contextual information, we adopted LLM-as-a-judge framework (Gu et al., 2024). Specifically, we utilized GPT-4o (OpenAI, 2024), genimi-2.5-flash (Comanici et al., 2025) and human to perform the evaluation. A total of 100 samples were randomly selected for this analysis. The evaluation was conducted using a carefully designed prompt, as illustrated in Figure 10. The results of the experiment are presented in Figure 8. The analysis reveals that the summarized reasoning chains exhibit a high degree of relevance and faithfulness. Specifically, the summaries achieved relevance scores of 90%, 92%, and 92%, indicating that the majority of the generated content aligns well with the given question-answer context. Similarly, faithfulness scores of 87%, 89%, and 90% demonstrate that the summaries largely adhere to the factual information provided by the context, with minimal hallucination. Furthermore, the alignment between LLM evaluations and human evaluations indicates the reliability and rationality of LLM-based judge.
Figure 8. Evaluating Relevance and Faithfulness of Reasoning Chain.
5.10.2. Quantitative comparison
In this experiment, we evaluate the impact of reasoning quality on QA performance by comparing two approaches: our $\text{Reasoning}_{\text{Align-GRAG}}$, where both the question and answer are included during reasoning chain generation, and $\text{Reasoning}_{\text{Naive}}$, where only the question is used for reasoning chain generation. The results in Table 5 demonstrate that incorporating the answer during summary generation ($\text{Reasoning}_{\text{Align-GRAG}}$) consistently outperforms the naive approach ($\text{Reasoning}_{\text{Naive}}$) across all evaluation metrics of datasets. Specifically, for ExplaGraphs, the accuracy rises significantly from 0.8592 using a naive reasoning approach to 0.8992 when employing $\text{Reasoning}_{\text{Align-GRAG}}$. For SceneGraphs, the accuracy improves from 0.8404 to 0.8804, illustrating the advantage of our method in enhancing graph-based reasoning accuracy. For WebQSP, The F1 score of $\text{Reasoning}_{\text{Align-GRAG}}$ is 0.5445, which is higher than the 0.5209 achieved by $\text{Reasoning}_{\text{Naive}}$, indicating that our strategy improves the balance of precision and recall in QA tasks. Similarly, the Hit@1 value increases from 0.7444 to 0.7626, showing that our approach helps the model identify correct nodes more accurately in graph alignment. These findings confirm that our method can consistently generate high-quality reasoning chains, which in turn lead to significant performance improvements.
Table 5. Quantitative Comparison of Different Reasoning Quality.
Method ExplaGraphs SceneGraphs WebQSP
[…]
$\text{Reasoning}_{\text{Align-GRAG}}$ 0.8992 0.8804 0.5445 0.7626 0.5700
$\text{Reasoning}_{\text{Naive}}$ 0.8592 0.8404 0.5209 0.7444 0.5409
@@ 5.10.2. Quantitative comparison
Table 6. The case for summarizing the reasoning chain. The text from SUMMARIZE_PROMPT, Question, Answer, and Graph DataBase are concatenated as input to generate the Response from LLM. We can see that LLM’s summarization accurately identified middle and critical node in graph, ‘J.K. Rowling’, for answer.
@@
5.11. Case study
In this case study, we investigate how summarization enhances the discovery of critical intermediate nodes within reasoning chains, as illustrated in Table reftab:summarization. Using well-crafted prompts in conjunction with a graph database, we show that concise, targeted summaries can surface pivotal entities—such as ‘J.K. Rowling’—that bridge the semantic gap between an initial query and the final answer. Our analysis evaluates the extent to which summarization reliably extracts these key connectors, thereby enabling coherent multi-hop reasoning that would otherwise be obscured within the raw graph structure. We further demonstrate that, without the interpretive layer provided by summarization, the underlying graph database lacks the necessary contextual signals to assemble the correct reasoning path end-to-end, underscoring the value of summarization as a catalyst for interpretable and effective knowledge retrieval.
5.12. LLM Prompts
We introduce three prompts: the Summarization Prompt, the Generator Prompt, and the Relevance and Faithfulness Prompt. Each serves a specific function in enhancing the performance and evaluation of LLMs. Below is a detailed explanation of their purposes and applications, as illustrated in Figure 9, Figure 11, and Figure 10.
@@ 5.12. LLM Prompts
Figure 9. Prompt for Summarization.
The Summarization Prompt is specifically designed to generate a comprehensive and cohesive summary by synthesizing three key inputs: a given question, its corresponding answer, and the relevant textualized graph data. The goal of this prompt is to create a logical connection between the question and the answer, while also integrating information extracted from the provided graph data. This ensures the output not only answers the question but also demonstrates an understanding of the contextual relationships within the graph. By combining these elements, the prompt facilitates the generation of a rich, informative summary that serves as a bridge between structured and unstructured data.
Figure 10. Evaluation of Relevance and Faithfulness Prompt.
@@
Figure 11. Prompt for Generator.
The Relevance and Faithfulness Prompt is an evaluation-focused tool that serves two distinct but complementary purposes:
  • Relevance Evaluation: This component assesses whether the reasoning chain provided is relevant to the given question. A relevant reasoning chain must contain information that contributes meaningfully to answering the question, even if the contribution is partial. The model is required to return one of two labels: Relevant or Irrelevant, with no additional commentary.
  • Faithfulness Evaluation: This component measures the faithfulness of the reasoning chain in responding to the question. A faithful reasoning chain should accurately represent the information in the context, avoid fabrications, and not contradict the provided data. Similarly, the model outputs one of two labels: Faithful or Not Faithful. This dual evaluation ensures the generated responses are both accurate and contextually grounded.
The Generator Prompt is tailored to the dataset being used, with its structure varying depending on the specific requirements of the dataset. For WebQSP and SceneGraphs, the prompt focuses on answering a question by leveraging the provided textualized graph data. This encourages the model to extract and synthesize relevant information from graph representations to deliver a concise and accurate response. For ExplaGraphs, the prompt shifts its focus to reasoning. It requires the model to evaluate the relationship between two arguments (e.g., whether they support or counter each other), based on the provided textualized graph data. The model outputs a single-word response, either support or counter, reflecting the nature of the relationship.
6. Conclusion
This work introduces Align-GRAG, a novel framework designed to address the core challenges in graph-based retrieval-augmented generation: the inclusion of irrelevant knowledge during retrieval and the representation gap between graph embeddings and language model representations. To tackle these issues, Align-GRAG incorporates an innovative graph aligner module that performs dual alignment. The Graph Node Alignment mechanism prunes irrelevant nodes, preserving critical structural information, while Graph Representation Alignment bridges the semantic gap, ensuring a unified representation space between graph and language models. Extensive experiments on the GraphQA benchmark demonstrate that Align-GRAG consistently outperforms existing baselines across multiple tasks, such as commonsense reasoning and knowledge graph understanding. By effectively integrating graph reasoning with LLM-based generation, Align-GRAG achieves more accurate, context-aware, and reliable outputs. This work paves the way for future advancements in GRAG systems, offering a robust solution for leveraging structured graph data in retrieval-augmented generation tasks.
7. Limitations